OpenStudy (anonymous):

write and equation for the line that is perpendicular to the line y=4x+5 and that passes through the point (0,-3)

OpenStudy (anonymous):

remember that a line perpendicular would have the slope of the opposite signed reciprocal of this line. and you would plug your point in to point slope form with your slope to get the equation y-y=m(x-x) point slope

OpenStudy (anonymous):

Thank you! (:

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

WAAAAIIT. LOL im confused cause i only have 1 point and not two..

OpenStudy (anonymous):

think of your other point as a variable in the equation y=mx+b where m is the slope and b is the y intercept.

OpenStudy (anonymous):

put your point and slope into point slope form and solve for y

OpenStudy (anonymous):

so i just put some random number into the equation?

OpenStudy (anonymous):

no you have a point to plug in. (0,-3) and m=-1/4 so you have enough to be left with 2 variables

OpenStudy (anonymous):

so its -3=-1/4(0)+5?

OpenStudy (anonymous):

no \[y+3=-1/4(x-0)\]

OpenStudy (anonymous):

your old y intercept isn't related to this

OpenStudy (anonymous):

ohhhh okay

OpenStudy (anonymous):

sorry my algebra teacher didn't teach us this yet =______= an just assigned us the work

OpenStudy (anonymous):

and what do i do with the my old y-intercept?

OpenStudy (anonymous):

its irrelevant to the new lines equation. graph your first line. then graph your second line and you will see what im talking about the new line will have its own y intercept

OpenStudy (anonymous):

what did you get. there is only 2 steps to this problem once in point slope

OpenStudy (anonymous):

i got y=-1/4x-3

OpenStudy (anonymous):

perfect did you graph them. even if it doesn't ask you to graphing them will help your intuitive thinking.

OpenStudy (anonymous):

yeah '

OpenStudy (anonymous):

Thank you! (: haha i appreciate it.